OpenStudy (anonymous):

Simplify the expression the root of negative four all over the quantity of five plus two i minus the quantity of three minus four i.

OpenStudy (anonymous):

\[\frac{\sqrt{-4}}{(5+2i)-(3-4i)}\]

OpenStudy (anonymous):

@campbell_st think you can help?

OpenStudy (campbell_st):

well you can simplify the denominator as \[5 +2i - 3 + 4i = 2 + 7i\] the numerator is \[\sqrt{-4} = \sqrt{4i^2} = \pm 2i\] you probably only need to use 2i so the problem is now \[\frac{2i}{2 + 7i}\] you'll probably need to rationalise the denominator... to get the final solution. hope it helps

OpenStudy (anonymous):

7i? i thought it was 6i .@campbell_st

OpenStudy (anonymous):

the answers they give are 3+i/10 3-i/10 3+1/8 3-i/8 @campbell_st

OpenStudy (anonymous):

@ganeshie8 think you can help?

OpenStudy (anonymous):

@Hero @whpalmer4 how about yall?

OpenStudy (campbell_st):

oops yes it is 6i... so to rationalise the denominator use the idea of the difference of 2 squares \[\frac{2i}{2 + 6i} \times \frac{2 - 6i}{2-6i}\] which becomes \[\frac{2i(2 - 6i)}{4 + 36}\] can you finish it from here...?

OpenStudy (anonymous):

\[\frac{ 4i-12i^2 }{ 40 }\]

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well remember \[i^2 = -1\] so its \[\frac{4i + 12}{40}\] remove the commone factor of 4 for the answer